# MR. Guard: Multilingual Reasoning Guardrail using Curriculum Learning

### v1

@@ 4.2 Multilingual Content Moderation
Benchmark: Our experiments cover 5 recent multilingual safety benchmarks: PTP_wildchat (Jain et al., 2024) [^5] We only select wildchat subset. (wildchat), RTP_LX (de Wynter et al., 2024), aya-red-teaming (Aya)(Aakanksha et al., 2024) , MultiJail (Deng et al., 2023), and Xsafety (Wang et al., 2023). We define five in-domain languages—English (EN), Arabic (AR), Spanish (ES), Chinese (ZH), and Russian (RU), which are included in the training data. To further assess generalization, we also evaluate on three out-of-domain languages (listed in Table 6) that are not included in the training set but are in the test datasets. Details of the evaluation benchmark are provided in Appendix.
@@
Figure 4: F1 score breakdown on the XSafety dataset, evaluated across 8 target languages. Here EN, AR, ES, RU, and ZH-Hans are in-domain languages, and FR, HI, JA are out-of-domain languages.
Baselines We compare our guardrail against several recent content moderation guardrails, both with and without reasoning capabilities. The configurations and details of the different baseline models are summarized in Table 1.
Table 2 summarizes the performance of our model on multilingual content moderation across several benchmark datasets. We observe that training on English-only data negatively impacts generalization to non-English languages. For example, although Aegis-2.0 and WildGuard perform well on English inputs, its performance degrades significantly on non-English data compared to LLaMA-Guard-3, which is trained with broader multilingual coverage. Moreover, models with reasoning capabilities demonstrate stronger generalization across datasets and languages, particularly in capturing language-specific nuances, as the test datasets are naturally generated or annotated by native speakers. Notably, our model achieves SOTA performance using a significantly smaller training set than other baselines as shown in Table 1, highlighting the sample efficiency of our approach for content moderation. In addition, the comparison between models trained with and without curriculum learning shows that gradually increasing the difficulty of training inputs, based on linguistic and cultural complexity, further enhances the model’s multilingual understanding. This finding underscores the value of curriculum-based learning strategies in improving robustness and generalization for multilingual safety tasks. Additionally, Figure 5 presents example reasoning generated by our multilingual guardrail, helping users understand the rationale behind its safety decisions.
@@
Moreover, we present a language-wise breakdown in Figure 3 and Figure 4. Our model consistently outperforms the baselines across a wide range of languages, although some baselines achieve comparable results on the English subset. For parallel datasets such as XSafety (Figure 4), our model exhibits a much smaller performance drop between ID and OOD languages compared to other baselines. This further demonstrates the robustness and cross-lingual generalization capabilities of our model.
4.3 Robustness to Multilingual Variations
In this section, we investigate the potential of using guardrails to identify unsafe prompts that involve variations specific to the multilingual setting. We consider two existing multilingual attacks: 1) Yoo et al. (2024) generates code-switching prompts using parallel datasets and GPT (CSRT); 2) Sandwich attack (Upadhayay and Behzadan, 2024) (Sandwich), where jailbreaking prompts are embedded within benign prompts in lower-resource languages. Both attack strategies have demonstrated that LLMs are more vulnerable to these challenging variants and are more likely to produce harmful responses. We crafted the variations of MultiJail and XSafety dataset using those two adversarial attacks, and examples are shown in Figure 8666The configuration and details of the generated attack are in D.
We benchmark several guardrail methods against adversarial multilingual attacks. Table 3 and 4 reports the F1 scores before and after the attacks, along with the corresponding changes. As shown in the table, all methods experience a decline in F1 score after attack, demonstrating the effectiveness of both adversarial strategies. Notably, our method not only outperforms the baselines but also exhibits a smaller reduction in F1 score. Our experiments show that incorporating reasoning alongside safety classification significantly enhances the guardrail’s robustness against multilingual adversarial prompts.
@@ 4.3 Robustness to Multilingual Variations
Ours 98.22 96.68 1.54
(a) MultiJail Dataset
Models EN $\uparrow$ Avg-CSRT $\uparrow$ $\Delta$downarrow
DUO-Guard 75.71 70.19 5.52
GuardR 85.99 84.42 1.57
LlaMa-Guard-3 64.48 61.08 3.40
Aegis-2.0 66.22 39.30 26.92
Wildguard 91.40 84.40 7.00
Ours 92.71 91.51 1.21
(b) XSafety Dataset
Table 3: F1 scores on code-switching prompts evaluated on the MultiJail and XSafety datasets. The best-performing results across models are highlighted in bold. $\Delta$ represents the difference between the F1 score on English prompts and the averaged F1 score over all code-switching variants across both ID and OOD languages.
@@
Ours 96.50 92.37 4.13
(a) MultiJail Dataset
Models
Avg-
Orig
$\uparrow$
Avg-
Sandwich
$\uparrow$
$\Delta$$\downarrow$
DUO-Guard 64.77 1.23 63.53
GuardR 82.25 74.83 7.42
LlaMa-Guard-3 61.50 6.37 55.13
Aegis-2.0 35.32 1.12 34.20
Wildguard 69.21 42.38 26.83
Ours 93.48 81.13 12.36
(b) XSafety Dataset
Table 4: F1 scores on sandwich attacks evaluated on the MultiJail and XSafety datasets. The best-performing results across models are highlighted in bold. Avg-Orig indicates the average F1 score on before attack, and the average F1 score after sandwich attack across both ID and OOD languages. $\Delta$ represents the difference between them.
5 Ablation Study
In this section, we conduct an ablation study to investigate the components of the reward function to show that all of them help improve the generalization of our guardrail’s performance on different languages.
@@ 5 Ablation Study
Table 5: F1 scores for the ablation study. Curr denotes GRPO with curriculum learning. $\mathcal{R}_{f+a}$ represents the combination of the format reward and accuracy reward. $\mathcal{R}_{u}$ corresponds to the uncertainty reward, and $\mathcal{R}_{l}$ denotes the language reward.
Based on the results in Table 5, we observe that both curriculum learning and uncertainty are essential for improving the guardrail’s performance in identifying safe and unsafe prompts. Although removing the language reward leads to slightly better performance across different datasets, we find that the resulting model predominantly generates English reasoning. In practice, however, it is important for the guardrail to produce reasoning in the corresponding input language, making multilingual reasoning generation a valuable capability despite the marginal trade-off in accuracy.
@@ 6 Conclusion
In this work, we propose a novel approach for building a guardrail for multilingual prompt moderation. Our method consists of three key stages: synthetic data generation, supervised fine-tuning, and reinforcement learning, where we adopt GRPO under a multi-stage curriculum setting. We conduct comprehensive experiments across multiple, diverse, and realistic multilingual content moderation benchmarks, demonstrating that our guardrail achieves SOTA performance with reasoning. The reasoning ability enables the multilingual users to understand the decision from the guardrail. We believe this work would be an important step toward enhancing the safety of utilizing LLMs in a multilingual world.
@@ Appendix B Experiment Setup
We used the Huggingface framework (Wolf et al., 2020) to load dataset and evaluate the guardrails and applied the default greedy decoding for decoder-based guardrails. Our training is performed on 4 NVIDIA RTX A100 (80G) GPUs, and vLLM (Kwon et al., 2023) is used to optimize inference speed.
For training data configuration, during the SFT stage, we train on the full English seed dataset (20.3K examples) combined with the translations generated from the sampled 2,000 seed examples. At the GRPO stage, we subsample another 2,000 seed examples and generate the challenging variants for curriculum learning, and we additionally include the seed English samples in the first curriculum stage to avoid losing its English ability as described in Section 3.3.
We use the TRL library (von Werra et al., 2020) for both SFT and GRPO stage. For both stages, we set the LoRA rank and alpha to 32, with a dropout rate of 0.1. During SFT, we use a learning rate of SFT is $52e-$ 5 and train for 3 epochs. For GRPO, we set the learning rate to $51e-$ 5 and the number of training epoch is 1. We conduct a hyperparameter sweep over LoRA rank and alpha values $\{8,16,32\}$, and select the best configuration based on performance on the Aegis-2.0 validation dataset.
@@ Appendix D Multilingual Variants
Appendix E Additional Results

### v2

@@ 4.2 Multilingual Content Moderation
Benchmark:
Our experiments cover 5 recent multilingual safety benchmarks: PTP_wildchat (Jain et al., 2024) [^5] We only select wildchat subset. (Wildchat), RTP_LX (de Wynter et al., 2024), aya-red-teaming (Aya)(Aakanksha et al., 2024) , MultiJail (Deng et al., 2023), and XSafety (Wang et al., 2023). We define five in-domain languages—English (EN), Arabic (AR), Spanish (ES), Chinese (ZH), and Russian (RU), which are included in the training data. To further assess generalization, we also evaluate on three out-of-domain languages (listed in Table 7) that are not included in the training set but are in the test datasets. Details of the evaluation benchmark are provided in Appendix C.
@@
Baselines:
We compare our guardrail against several recent content moderation guardrails, both with and without reasoning capabilities. The configurations and details of the different baseline models are summarized in Table 1.
Table 2 summarizes MrGuard’s performance on several multilingual moderation benchmarks. Across both in-domain and out-of-domain languages, MrGuard consistently outperforms all baselines by a substantial margin. Additionally, we observe that guardrails with reasoning capabilities (GuardR and ours) generalize better across datasets and languages, but with our approach, MrGuard achieves state-of-the-art performance in multilingual scenarios. This showcases that MrGuard effectively captures language-specific nuances, as the test datasets are naturally generated or annotated by native speakers. Note that we restrict our experiments to compact models ($\leq$ 7 B parameters) to ensure low latency and easy deployment, in line with other guardrails (see Table 1). Even at this scale, adding reasoning at the post-training stage yields large gains in multilingual classification.
Moreover, we present a language-wise breakdown in Figure 3 with more results in Appendix B. Our model consistently outperforms the baselines across a wide range of languages, although some baselines achieve comparable results on the English subset. We also observe that training on English-only data negatively impacts generalization to non-English languages. For example, although Aegis-2.0 and WildGuard perform well on English inputs, their performance degrades significantly on non-English data, whereas MrGuard maintains high performance and shows minimal performance drop, even on unseen languages. This further demonstrates the robustness and cross-lingual generalization capabilities of our model.
Furthermore, Figure 4 presents example reasoning generated by our multilingual guardrail in various languages, showing that the reasoning accurately analyzes and justifies the safety prediction of the input prompts. The reasoning can help users understand the rationale of MrGuard behind its safety decisions.
@@
4.3 Robustness to Multilingual Perturbations
In this section, we investigate the potential of using guardrails to identify unsafe prompts that involve perturbations specific to the multilingual setting. We consider two existing multilingual attacks: 1) Yoo et al. (2024) generates code-switching prompts using two parallel datasets, MultiJail and XSafety, and GPT (CSRT) 2) Sandwich attack (Upadhayay and Behzadan, 2024) (Sandwich), where jailbreaking prompts are embedded within benign prompts in lower-resource languages. Both attack strategies have demonstrated that LLMs are more vulnerable to these challenging input variants and are more likely to produce harmful responses. We crafted the variations of MultiJail and XSafety datasets using those two adversarial attacks, and examples are shown in Figure 9666The configuration and details of the generated attack are in Appendix D..
We benchmark several guardrail methods against the adversarial multilingual attacks. Table 3 and 4 reports the F1 scores before and after the attacks, along with the corresponding performance changes. As shown in the tables, all methods experience a decline in F1 score after the attack, demonstrating the effectiveness of both adversarial strategies. Notably, our method not only outperforms the baselines but also exhibits a smaller reduction in F1 score. Our experiments show that incorporating reasoning alongside safety classification significantly enhances the guardrail’s robustness against multilingual adversarial prompts.
@@ 4.3 Robustness to Multilingual Perturbations
MrGuard (Ours) 98.22 96.68 1.54
Table 3: F1 scores on code-switching prompts evaluated on the MultiJail datasets. The best-performing results across models are highlighted in bold. $\Delta$ represents the difference between the F1 score on English prompts and the averaged F1 score over all code-switching variants across both ID and OOD languages.
@@
MrGuard (Ours) 96.50 92.37 4.13
Table 4: F1 scores on sandwich attacks evaluated on the MultiJail dataset. The best-performing results across models are highlighted in bold. Avg-Orig indicates the average F1 score on before attack, and the average F1 score after sandwich attack across both ID and OOD languages. $\Delta$ represents the difference between them.
5 Discussion
In this section, we conduct a deeper analysis of our framework and results, including ablation experiments of the proposed approach, evaluation of the fidelity of reasoning and safety predictions, and cross-language consistency.
5.1 Ablation Study
In this section, we conduct an ablation study to investigate the effectiveness of GRPO and curriculum learning, and various components of the reward function to show that all of them help improve the generalization of our guardrail’s performance on different languages across different datasets.
@@ 5.1 Ablation Study
Table 5: F1 scores for the ablation study. Curr denotes GRPO with curriculum learning. $\mathcal{R}_{f+a}$ represents the combination of the format reward and accuracy reward. $\mathcal{R}_{u}$ corresponds to the uncertainty reward, and $\mathcal{R}_{lang}$ denotes the language reward.
Based on the results in Table 5, we first observe that both GRPO and curriculum learning significantly improve the performance compared to $\pi_{\text{sft}}$. Consistent with prior work (DeepSeek-AI, 2025), post-training with GRPO improves the generalization across different datasets and enhances reasoning abilities. Moreover, the comparison between models trained with and without curriculum learning shows that gradually increasing the difficulty of training inputs, based on linguistic and cultural complexity, further enhances the model’s multilingual understanding. This finding underscores the value of curriculum-based learning strategies in improving robustness and generalization for multilingual safety tasks.
Furthermore, we show that all components in the reward functions positively contributes to the overall performance of MrGuard. Although removing the language reward leads to slightly better performance across different datasets, we find that the resulting model predominantly generates English reasoning. In practice, however, it is important for the guardrail to produce reasoning in the corresponding input language, making multilingual reasoning generation a valuable capability despite the marginal trade-off in accuracy. We leave the theoretical analysis of curriculum learning and reward function design of GRPO to future work.
5.2 Cross-lingual Consistency
One important characteristic for guardrails is that it assigns the same safety label to semantically equivalent prompts in different languages. To quantify this, we define the Cross-Lingual Consistency score as the fraction of parallel examples in which the model’s safety predictions agree across languages (She et al., 2024). We report consistency score on XSafety, a parallel dataset, comparing English with each target language across several models in Figure 5. From the results we observe that although our algorithm does not explicitly train for consistency, we still see improved consistency, especially for unsafe prompt classification. As shown in Figure 10, MrGuard exhibits a much smaller performance drop between ID and OOD languages compared to the other baselines.
Figure 5: Consistency score between English and non-English on the XSafety dataset, evaluated across 8 target languages. Here AR, ES, RU, and ZH are in-domain languages, and FR, HI, JA are out-of-domain languages. Higher the score the better.
5.3 Quality of Reasoning
LLMs are likely to produce hallucinations, even when guided via chain-of-thought. MrGuard is intended to help users and regulators inspect and trust its decisions, making fidelity measurement crucial. To this end, we employ a stronger LLM, GPT-4.1-mini, as a judge to automatically assess whether each explanation faithfully reflects the input and correctly drives the safety prediction. We define the Explanation Fidelity (EF) score as the fraction of reasoning sentences the judge labels as coherent out of the total sentences, and the results are shown in Table 6. Moreover, it is important to maintain the reasoning language in the same language as the prompt. We report the Language Match (LM) rate, which captures the percentage of cases where the generated reasoning is in the same language as the prompt. The results show that our reasoning is aligned with the semantics and the language of the prompt.
Lang EN AR ZH RU HI
EF 87.39 80.57 93.33 86.53 88.97
LM 97.30 98.76 99.52 98.86 99.91
Table 6: Here we report two metrics on RTP_LX dataset to evaluate the quality of the reasoning, where EF is the explanation fidelity rate and LM indicates the language matching rate. Higher score the better.
@@ 6 Conclusion
In this work, we introduce MrGuard: a multilingual reasoning-enhanced guardrail for multilingual prompt moderation. Our method consists of three key stages: synthetic data generation, supervised fine-tuning, and reinforcement learning, where we adopt GRPO with a multi-stage curriculum that progressively introduces more cultural and language-specific elements. We conduct comprehensive experiments across multiple diverse and realistic multilingual content moderation benchmarks, including challenging scenarios involving code-switching, and demonstrate that our guardrail achieves state-of-the-art performance with reasoning. We also analyze the generated reasoning to validate its reliability and ensure consistent safety preservation across languages. The reasoning ability enables multilingual users to understand the decision from MrGuard. We believe this work is an important step toward enhancing the safety of LLMs in a multilingual world.
Limitations
Language and resource coverage Due to budget and computational limits, we generated synthetic data only for high- and mid-resource languages, and relied on Aegis-2.0 as our English seed dataset. Expanding to additional seed datasets and low-resource languages could further enhance model performance and broaden the safety taxonomy to better reflect diverse user needs. Additionally, while our guardrail demonstrates strong results on both in-distribution and out-of-distribution dataset and languages, the languages represented in our evaluation remain limited.
Potential Bias We use a single LLM as a judge to verify safety labels of translations, which may introduce bias inherent to the LLM. Additionally, our current evaluation of reasoning coherence and faithfulness between explanations and final safety predictions relies on automated heuristics, which may not perfectly align with human judgments.
Ethical Statement
Our works aims to improve LLM safety for multilingual users by introducing a multilingual reasoning guardrail, which is important for building a universally reliable LLM for safety-critical applications. The generated synthetic data and models will be released, accompanied by detailed usage guidelines to prevent misuse.
@@ Appendix B Experiment Setup
We used the Huggingface framework (Wolf et al., 2020) to load dataset and evaluate the guardrails and applied the default greedy decoding for all guardrails. Our training is performed on 4 NVIDIA RTX A100 (80G) GPUs, and vLLM (Kwon et al., 2023) is used to optimize inference speed.
For training data configuration, during the SFT stage, we train on the full English seed dataset (30.3K examples) combined with the translations generated from the sampled 2,000 seed examples. At the GRPO stage, we subsample another 2,000 seed examples and generate the challenging variants for curriculum learning, and we additionally include the seed English samples in the first curriculum stage to avoid losing its English ability as described in Section 3.3.
We use the TRL library (von Werra et al., 2020) for both SFT and GRPO stage. For both stages, we set the LoRA rank and alpha to 32, with a dropout rate of 0.1. During SFT, we use a learning rate of SFT is $52e-$ 5 and train for 3 epochs. For GRPO, we set the learning rate to $51e-$ 5 and the number of training epoch is 1. We conduct a hyperparameter sweep over LoRA rank and alpha values $\{8,16,32\}$, and select the best configuration based on performance on the Aegis-2.0 validation dataset. The results are the average of three training runs. All use of the packages and artifacts are consistent with their intended use and license. We used ChatGPT to refine short sentences and paragraphs and to check for grammar errors.
@@ C.1 Generation configuration
C.2 Human evaluation
We validated our LLM-generated synthetic translations (Section 3) by comparing the model’s safety judgments with human judgments. We have subsampled 100 sentences for ZH, RU, and AR. For each language, we have one volunteer, and the volunteers are native speakers in the target languages and are proficient in English. The guidelines of human evaluation is shown in Figure 8, and the results are listed in Table 9. The safety preservation rate is a three-point scale measurement (0-2). 0: The safety label is not preserved; 1: the safety label is preserved, but meaning is altered; 2: both safety label and original meaning are preserved.
Lang AR ZH RU
Translation 1.98 1.75 1.9
Local 1.67 1.73 1.91
Challenging 1.32 1.58 1.75
Table 9: Here we report human annotation on safety preservation rate of synthetic data generation. Higher score the better. The scale is from 0 to 2.
Figure 8: Guidelines for human annotation of synthetic multilingual training data.
@@ Appendix D Multilingual Variants
Appendix E GRPO algorithm
The objective function of GRPO is defined as follow. Let $\pi_{\theta}$ denote the language model parameterized by $\theta$, and $\pi_{\text{old}}$ be the model from a previous iteration. Given a prompt $p$, we sample a group of generations $\{o_{1},\ldots,o_{G}\}\sim\pi_{\text{old}}(\cdot\mid p)$, each associated with a reward $\{r_{1},\ldots,r_{G}\}$. Let $\mathbb{D}_{\text{KL}}$ denote the KL-divergence between two distributions. GRPO estimates the advantage of a generation $o_{i}$ using:
The GRPO objective is then defined as:
$-\beta\mathbb{D}_{\rm KL}[\pi_{\theta}||\pi_{\rm ref}].$
Appendix F Additional Results on QWEN
Table 10 shows additional results from training QWEN2.5-3B (Qwen et al., 2025) with our framework.
F1 RTP_LX Aya XSafety MultiJail
QWEN 89.76 98.26 95.43 91.56
Table 10: Here we report F1 scores of QWEN-2.5-3B across different datasets. We take the average across both in-domain and out-of-domain languages.
Appendix G Additional Results
@@
Figure 10: F1 score breakdown on the XSafety dataset, evaluated across 8 target languages.
@@
Figure 13: F1 score breakdown on the Multijail dataset, evaluated across 8 target languages.
We additionally perform code-switching and sandwich attack on XSafety dataset and the results are shown in Table 11 and Table 12.
Models EN $\uparrow$ Avg-CSRT $\uparrow$ $\Delta$downarrow
DUO-Guard 75.71 70.19 5.52
GuardR 85.99 84.42 1.57
LlaMa-Guard-3 64.48 61.08 3.40
Aegis-2.0 66.22 39.30 26.92
Wildguard 91.40 84.40 7.00
MrGuard 92.71 91.51 1.21
Table 11: F1 scores on code-switching prompts evaluated on the XSafety datasets. The best-performing results across models are highlighted in bold. $\Delta$ represents the difference between the F1 score on English prompts and the averaged F1 score over all code-switching variants across both ID and OOD languages.
Models
Avg-
Orig
$\uparrow$
Avg-
Sandwich
$\uparrow$
$\Delta$$\downarrow$
DUO-Guard 64.77 1.23 63.53
GuardR 82.25 74.83 7.42
LlaMa-Guard-3 61.50 6.37 55.13
Aegis-2.0 35.32 1.12 34.20
Wildguard 69.21 42.38 26.83
MrGuard 93.48 81.13 12.36
Table 12: F1 scores on sandwich attacks evaluated on the XSafety datasets. The best-performing results across models are highlighted in bold. Avg-Orig indicates the average F1 score on before attack, and the average F1 score after sandwich attack across both ID and OOD languages. $\Delta$ represents the difference between them.